OpenStudy (graciek):

Solve each problem. 2/3 = 1.2/x A. X=18 B. X=1.8 C. X=0.2 D. X=10 4/5 = 8/ x+2 A. X=4 B. X=9.5 C. X=2 D. X=8

OpenStudy (rushwr):

\[\frac{ 2 }{ 3 } =\frac{ 1.2 }{ x }\] \[x = \frac{ 1.2 * 3 }{ 2} = 1.8\]

OpenStudy (rushwr):

\[\frac{ 4 }{ 5 } = \frac{ 8 }{ (x+2)}\] \[4(x+2) = 8*5\] \[x+ 2 = \frac{ 8 *5 }{ 4 }= 10\] \[x= 8\]

OpenStudy (graciek):

Thanks!

OpenStudy (rushwr):

NO problem ! I hope u understand !